A train leaves a city heading west and travels at 50 miles per hour. Three hours later, a second train leaves fromt he same place and travels in the same directions at 65 miles per hour. how long will it take for the second train to overtake the first train?
Answer by kapilsinghi(68)

let both trains meet each other x hours after second train starts
Time taken by train 1 will be = x+3
distance covered by train 1 = 50(x+3)
Time taken by train 2 will be = x
distance covered by train 2 = 65(x)

bothe distances will be equal
so 65x = 50(x+3)
65x = 50x+150
65x - 50x = 150
15x = 150
x = 150/15
x = 10
ans 10 Hours
kapilsinghi123@gmail.com